Mathematics 242 – Principles of Analysis Definitions Quiz 3 – February 14, 2014

Your Name:

A) (5) Define: The real number a is lub(A) for a subset A of the real numbers.

B) (5) Give an example of a bounded set A of real numbers for which glb(A) ∈ A, but lub(A) ∉ A.